# Multi-CATE: Multi-Accurate Conditional Average Treatment Effect Estimation Robust to Unknown Covariate Shifts

### v1

@@ Results
Train Shift CForest S-learner DR-learner T-learner CForest S-learner DR-l T-learner
size degree OS OS OS Ridge Tree OS Ridge Tree RCT wRCT RCT wRCT RCT RCT wRCT
0 46.53 61.3 111.73 82.85 83.54 108.92 45.41 51.38 5.22 5.28 18.27 18.27 7.07 16.15 7.33
0.25 45.23 58.48 107.7 105.93 123.73 105.4 42.59 45.13 6.91 8.58 18.42 19.6 9.11 14.66 8.92
0.5 47.1 59.48 104.44 81.08 100.4 108.1 36.17 35.04 13.45 23.04 20.44 23.6 15.97 26.37 13.64
0.75 49.53 64.22 109.15 93.88 89.59 115.87 45.45 36.6 25.35 45.64 24.45 27.44 23.96 27.56 19.9
500 1 48.78 63.54 124.72 93.16 108.59 112.39 45.89 32.91 43.62 106.32 30.65 30.28 36 40.86 33.74
1.25 45.41 56.97 111.54 91.98 99.49 100.86 41.42 30.18 62.16 139.28 40.83 31.4 52.28 52.26 48.02
1.5 48.76 61.09 120.58 83.81 152.2 113.71 45.95 33.42 88.73 91.14 56.46 36.34 70.6 59.34 63.67
1.75 46.44 60.79 115.11 106.45 108.84 108.87 45.81 26.6 114.53 183.94 75.91 40.42 100.12 88.11 107.75
2 46.1 60.06 113.18 102.82 174.57 108.76 44.54 23.56 126.01 125.37 78.09 39.27 99.86 93.8 85.42
0 34.02 56.98 83.89 54.37 58.17 83.85 48.68 64.93 5.81 5.88 19.34 19.61 8.19 15.34 8.28
0.25 29.49 50.98 76.11 49.77 60.99 75.93 25.82 25.96 6.98 10.82 18.25 20.02 9.11 17.66 9.61
0.5 33.12 55.6 72.25 46.41 46.75 81.12 30.23 26.36 13.76 37.77 19.88 24.6 14.86 29.09 13.93
0.75 31.86 54.78 80.56 49.81 55.36 81.86 36.61 28.8 27.04 154.84 25.33 28.49 27.44 41.5 28.09
2000 1 33.65 55.77 72.6 46.45 60.35 80.97 38.76 23.86 44.3 63.07 31.02 31.36 37.28 39.17 32.74
1.25 33.01 56.3 87.1 50.7 64.66 84.99 43.36 23.91 70.32 67.38 46.2 33.52 61.94 67.42 56.73
1.5 33.69 55.48 77.48 51.87 70.75 81.06 43.36 22.2 88.01 93.85 58.31 34.63 74.66 66.4 71.68
1.75 31.71 54.91 79.32 50.4 76.92 82.71 42.06 19.67 111.74 147.67 68.42 35.19 89.55 65.77 82.72
2 32.33 56.07 67.16 51.21 58.54 82.06 41.47 19.17 129.68 111.65 77.32 38.03 100.22 131.54 79.72
0 28.05 51.13 72.69 38.9 48.91 72.77 28.94 31.78 5.44 5.67 18.39 18.49 8.02 15.26 7.95
0.25 28.66 51.48 70 38.2 51.76 74.17 38.98 32.72 7.07 11.5 18.86 21.25 9.13 16.56 9.78
0.5 27.74 50.57 66.08 38.05 47.27 72.25 26.5 24.19 14.89 46.8 20.9 27.29 16.06 22.96 16.65
0.75 29.41 52.81 79.39 44.81 49.55 75.39 33.09 23.02 26.21 103.1 24.51 28.51 24.75 41.25 21.19
3500 1 30.62 53.15 64.92 38.36 58.88 75 39.02 21.75 45.43 127.32 33.08 29.67 41.02 47.21 34.11
1.25 27.33 49.97 66.27 38.69 78.7 71.01 37.88 20.5 68.1 173.54 42.18 32.84 52.77 65.16 45.2
1.5 27 50.26 68.44 47.33 68.79 70.92 38.41 20.9 89.94 112.42 57.41 33.69 75.11 81.29 60.63
1.75 29.01 53.2 68.33 44.42 72.64 74.55 42.47 18.1 109.79 136.09 66.94 36.19 86.19 68.07 80.93
2 29.14 53.3 66.32 47.85 58.99 75.38 41.67 16.48 128.92 113.56 77.84 44.29 100.65 95.59 91.14
0 26.08 48.7 60.13 31.96 38.88 67.99 23.98 26.11 5.42 5.5 17.6 17.51 7.27 13.99 7.38
0.25 25.26 48.91 59.2 35.38 41.97 66.92 24.3 23.91 7.48 13.7 19.15 21.72 9.84 18.6 10.69
0.5 25.51 47.96 58.62 35.94 39.65 66.99 25.84 22.22 13.46 37.92 19.55 24.89 14.6 27.59 15.32
0.75 26.09 48.06 56.54 37.49 39.72 66.26 29.67 20.88 25.74 53.22 23.56 27.51 24.02 40.76 20.13
5000 1 26.2 48.63 60.2 35.77 39.01 67.02 35.05 20.49 44.87 100.33 31.93 30.69 38.84 46.02 38.92
1.25 24.7 47.38 64.71 42.64 57.49 65.43 37.81 19.17 67.81 108.59 43.38 33.45 54.84 61.97 50.91
1.5 26.65 50.79 63.56 43.44 54.69 69.8 42.46 17.81 94.26 124.26 59.78 35.38 78.68 69.38 89.64
1.75 28.42 50.98 62.66 50.51 59.41 69.61 42.37 16.1 110.83 190.54 68.43 37.49 89.63 90.15 88.29
2 25.92 49.49 58.89 41.56 61.87 67.18 40.06 15.78 126.05 90.74 74.91 35.83 98.03 85.04 77.14
@@ D.3 WHI Data Application
We consider a case study using clinical trial and observational data from the Women’s Health Initiative [Machens and Schmidt‐Gollwitzer, 2003]. A focus of this study was to investigate the effectiveness of hormone replacement therapy (HRT) treatment in preventing the onset of chronic (cardiovascular) diseases. As the observational study and clinical trial data led to conflicting findings, the WHI study has become a prime example of how confounding in observational data can introduce bias and, in this case, suggest overly optimistic results (for more detail, see Kallus and Zhou 2018). In this setting, we study how multi-accurate CATE estimators that are “warm-started” with observational data and have access to small samples from the clinical trial compare to estimators that draw on either observational or clinical trial data only.
@@
We estimate DR-learner and T-learner using multi-calibration boosting with samples of clinical trial data. The MCBoost hyperparameter settings are shown in Table 12b.
@@ Results
Figure 15: MSE by ’true’ CATE estimation method and clinical trial training set size with extended set of covariates (WHI Data Application). The distribution of MSE scores over sampling repetitions is plotted. Multicalibration boosting yields little improvement as CForest-OS and T-learner-OS already achieve low MSE.

### v2

@@ Results
Train Shift KL CForest S-learner DR-learner T-learner CForest S-learner DR-l T-learner
size degree Div. OS OS OS Ridge Tree OS Ridge Tree RCT wRCT RCT wRCT RCT RCT wRCT
0 2.98 46.53 61.3 111.73 82.85 83.54 108.92 45.41 51.38 5.22 5.28 18.27 18.27 7.07 16.15 7.33
0.25 3.19 45.23 58.48 107.7 105.93 123.73 105.4 42.59 45.13 6.91 8.58 18.42 19.6 9.11 14.66 8.92
0.5 3.62 47.1 59.48 104.44 81.08 100.4 108.1 36.17 35.04 13.45 23.04 20.44 23.6 15.97 26.37 13.64
0.75 4.47 49.53 64.22 109.15 93.88 89.59 115.87 45.45 36.6 25.35 45.64 24.45 27.44 23.96 27.56 19.9
500 1 5.78 48.78 63.54 124.72 93.16 108.59 112.39 45.89 32.91 43.62 106.32 30.65 30.28 36 40.86 33.74
1.25 7.79 45.41 56.97 111.54 91.98 99.49 100.86 41.42 30.18 62.16 139.28 40.83 31.4 52.28 52.26 48.02
1.5 10.33 48.76 61.09 120.58 83.81 152.2 113.71 45.95 33.42 88.73 91.14 56.46 36.34 70.6 59.34 63.67
1.75 13.82 46.44 60.79 115.11 106.45 108.84 108.87 45.81 26.6 114.53 183.94 75.91 40.42 100.12 88.11 107.75
2 16.98 46.1 60.06 113.18 102.82 174.57 108.76 44.54 23.56 126.01 125.37 78.09 39.27 99.86 93.8 85.42
0 2.99 34.02 56.98 83.89 54.37 58.17 83.85 48.68 64.93 5.81 5.88 19.34 19.61 8.19 15.34 8.28
0.25 3.00 29.49 50.98 76.11 49.77 60.99 75.93 25.82 25.96 6.98 10.82 18.25 20.02 9.11 17.66 9.61
0.5 3.67 33.12 55.6 72.25 46.41 46.75 81.12 30.23 26.36 13.76 37.77 19.88 24.6 14.86 29.09 13.93
0.75 4.61 31.86 54.78 80.56 49.81 55.36 81.86 36.61 28.8 27.04 154.84 25.33 28.49 27.44 41.5 28.09
2000 1 5.99 33.65 55.77 72.6 46.45 60.35 80.97 38.76 23.86 44.3 63.07 31.02 31.36 37.28 39.17 32.74
1.25 7.48 33.01 56.3 87.1 50.7 64.66 84.99 43.36 23.91 70.32 67.38 46.2 33.52 61.94 67.42 56.73
1.5 10.33 33.69 55.48 77.48 51.87 70.75 81.06 43.36 22.2 88.01 93.85 58.31 34.63 74.66 66.4 71.68
1.75 13.73 31.71 54.91 79.32 50.4 76.92 82.71 42.06 19.67 111.74 147.67 68.42 35.19 89.55 65.77 82.72
2 17.68 32.33 56.07 67.16 51.21 58.54 82.06 41.47 19.17 129.68 111.65 77.32 38.03 100.22 131.54 79.72
0 2.99 28.05 51.13 72.69 38.9 48.91 72.77 28.94 31.78 5.44 5.67 18.39 18.49 8.02 15.26 7.95
0.25 2.97 28.66 51.48 70 38.2 51.76 74.17 38.98 32.72 7.07 11.5 18.86 21.25 9.13 16.56 9.78
0.5 3.50 27.74 50.57 66.08 38.05 47.27 72.25 26.5 24.19 14.89 46.8 20.9 27.29 16.06 22.96 16.65
0.75 4.45 29.41 52.81 79.39 44.81 49.55 75.39 33.09 23.02 26.21 103.1 24.51 28.51 24.75 41.25 21.19
3500 1 6.00 30.62 53.15 64.92 38.36 58.88 75 39.02 21.75 45.43 127.32 33.08 29.67 41.02 47.21 34.11
1.25 7.50 27.33 49.97 66.27 38.69 78.7 71.01 37.88 20.5 68.1 173.54 42.18 32.84 52.77 65.16 45.2
1.5 10.40 27 50.26 68.44 47.33 68.79 70.92 38.41 20.9 89.94 112.42 57.41 33.69 75.11 81.29 60.63
1.75 13.87 29.01 53.2 68.33 44.42 72.64 74.55 42.47 18.1 109.79 136.09 66.94 36.19 86.19 68.07 80.93
2 17.16 29.14 53.3 66.32 47.85 58.99 75.38 41.67 16.48 128.92 113.56 77.84 44.29 100.65 95.59 91.14
0 3.12 26.08 48.7 60.13 31.96 38.88 67.99 23.98 26.11 5.42 5.5 17.6 17.51 7.27 13.99 7.38
0.25 3.06 25.26 48.91 59.2 35.38 41.97 66.92 24.3 23.91 7.48 13.7 19.15 21.72 9.84 18.6 10.69
0.5 3.59 25.51 47.96 58.62 35.94 39.65 66.99 25.84 22.22 13.46 37.92 19.55 24.89 14.6 27.59 15.32
0.75 4.56 26.09 48.06 56.54 37.49 39.72 66.26 29.67 20.88 25.74 53.22 23.56 27.51 24.02 40.76 20.13
5000 1 5.82 26.2 48.63 60.2 35.77 39.01 67.02 35.05 20.49 44.87 100.33 31.93 30.69 38.84 46.02 38.92
1.25 7.69 24.7 47.38 64.71 42.64 57.49 65.43 37.81 19.17 67.81 108.59 43.38 33.45 54.84 61.97 50.91
1.5 10.31 26.65 50.79 63.56 43.44 54.69 69.8 42.46 17.81 94.26 124.26 59.78 35.38 78.68 69.38 89.64
1.75 13.72 28.42 50.98 62.66 50.51 59.41 69.61 42.37 16.1 110.83 190.54 68.43 37.49 89.63 90.15 88.29
2 17.32 25.92 49.49 58.89 41.56 61.87 67.18 40.06 15.78 126.05 90.74 74.91 35.83 98.03 85.04 77.14
@@ D.3 WHI Data Application
We consider a case study using clinical trial and observational data from the Women’s Health Initiative [Machens and Schmidt-Gollwitzer, 2003]. A focus of this study was to investigate the effectiveness of hormone replacement therapy (HRT) treatment in preventing the onset of chronic (cardiovascular) diseases. As the observational study and clinical trial data led to conflicting findings, the WHI study has become a prime example of how confounding in observational data can introduce bias and, in this case, suggest overly optimistic results (for more detail, see Kallus and Zhou 2018). In this setting, we study how multi-accurate CATE estimators that are “warm-started” with observational data and have access to small samples from the clinical trial compare to estimators that draw on either observational or clinical trial data only.
@@
We estimate DR-learner and T-learner using multi-calibration boosting with samples of clinical trial data. The MCBoost hyperparameter settings are shown in Table 13b. Hyperparameters of the baseline CATE learner are listed in Table 14.
@@
Table 14: Hyperparameter settings of (baseline) CATE learners. Default settings are used for parameters not listed.
Method Implementation Hyperparameter Value
CForest grf num.trees 2000
mtry sqrt(p)+20
sample.fraction 0.5
honesty.fraction 0.5
min.node.size 5
T-,DR-learner regression_forest num.trees 2000
mtry sqrt(p)+20
sample.fraction 0.5
honesty.fraction 0.5
min.node.size 5
S-learner ranger num.trees 500
mtry sqrt(p)
min.node.size 5
replace TRUE
sample.fraction 1
@@ Results
Figure 15: MSE by “true” CATE estimation method and clinical trial training set size with extended set of covariates (WHI Data Application). The distribution of MSE scores over sampling repetitions is plotted. Multicalibration boosting yields little improvement as CForest-OS and T-learner-OS already achieve low MSE.